# ACEBench: Who Wins the Match Point in Tool Learning?

### v2

@@ A.2 Special Example
Appendix B Evaluation Inference Prompts
B.1 Normal Prompt
@@ B.3 Agent Prompt
Appendix C Error Examples
C.1 Error Examples of Normal Data
@@
C.2 Error Examples of Special Data
@@
Appendix D Formula for Overall Accuracy
@@
Appendix E The Construction of Agent Data
Second, implement logical abstraction for each function through code, designing core processes such as user authentication, cost calculation, payment processing, and information recording, ensuring the scalability and robustness of the code logic while also addressing exception handling comprehensively.
Appendix F Detailed Experimental Results
Table 6: Accuracy Evaluation of Different Models on ACEBench (English)
@@ Appendix F Detailed Experimental Results
GPT-4o 0.90 0.78 0.68 0.80 0.78 0.83 0.93 0.56 0.82
GPT-4-Turbo 0.91 0.80 0.69 0.80 0.88 0.84 0.82 0.62 0.80
Qwen-Max 0.88 0.75 0.61 0.74 0.82 0.80 0.74 0.60 0.75
DeepSeek-Chat 0.88 0.78 0.63 0.76 0.78 0.80 0.73 0.34 0.71
GPT-4o-mini 0.84 0.73 0.59 0.74 0.72 0.76 0.77 0.28 0.69
Gemini-1.5-Pro 0.82 0.73 0.61 0.74 0.72 0.76 0.77 0.26 0.69
Claude-3-5-Sonnet 0.67 0.64 0.46 0.58 0.68 0.62 0.73 0.44 0.62
Doubao-Pro-32k 0.75 0.58 0.52 0.70 0.54 0.66 0.51 0.27 0.56
@@
Qwen2.5-Coder-32B-Instruct 0.86 0.73 0.59 0.76 0.72 0.77 0.80 0.50 0.74
Qwen2.5-72B-Instruct 0.81 0.74 0.64 0.76 0.80 0.77 0.74 0.38 0.70
Llama-3.1-70B-Instruct 0.84 0.72 0.61 0.74 0.66 0.76 0.29 0.41 0.58
Qwen2.5-7B-Instruct 0.70 0.57 0.49 0.62 0.58 0.63 0.49 0.15 0.52
Qwen2.5-Coder-7B-Instruct 0.73 0.64 0.52 0.70 0.58 0.67 0.25 0.18 0.48
DeepSeek-Coder-V2-Lite-Instruct 0.72 0.58 0.50 0.62 0.60 0.64 0.39 0.03 0.48
Watt-Tool-8B 0.85 0.72 0.57 0.70 0.62 0.75 0.02 0.01 0.44
Hammer2.1-7b 0.71 0.62 0.43 0.64 0.52 0.63 0.03 0.15 0.40
Phi-3-mini-128k-Instruct 0.66 0.49 0.31 0.58 0.32 0.54 0.12 0.00 0.34
Llama-3.1-8B-Instruct 0.51 0.49 0.28 0.60 0.56 0.48 0.15 0.07 0.33
xLAM-7b-r 0.62 0.42 0.32 0.66 0.00 0.49 0.04 0.10 0.31
Llama-3.2-3B-Instruct 0.32 0.22 0.09 0.34 0.32 0.26 0.09 0.00 0.17
Hammer2.1-3b 0.33 0.14 0.07 0.36 0.32 0.26 0.01 0.01 0.15
Table 7: Accuracy Evaluation of Different Models on ACEBench (Chinese)
@@
GPT-4o-2024 0.97 0.91 0.86 0.90 0.88 0.93 0.93 0.71 0.90
GPT-4-Turbo 0.96 0.89 0.86 0.92 0.84 0.92 0.91 0.72 0.89
Qwen-Max 0.94 0.86 0.76 0.92 0.84 0.89 0.74 0.69 0.82
DeepSeek-Chat 0.95 0.91 0.91 0.90 0.88 0.93 0.73 0.35 0.79
GPT-4o-mini 0.89 0.79 0.74 0.80 0.84 0.83 0.81 0.39 0.76
Claude-3-5-Sonnet 0.87 0.81 0.79 0.84 0.76 0.84 0.82 0.35 0.76
Gemini-1.5-Pro 0.87 0.80 0.68 0.86 0.84 0.82 0.80 0.25 0.73
Doubao-Pro-32k 0.84 0.53 0.64 0.82 0.78 0.75 0.59 0.24 0.63
@@
Qwen2.5-Coder-32B-Instruct 0.94 0.89 0.83 0.90 0.90 0.91 0.81 0.71 0.85
Qwen2.5-72B-Instruct 0.92 0.86 0.75 0.90 0.82 0.87 0.77 0.57 0.80
Llama-3.1-70B-Instruct 0.81 0.65 0.66 0.84 0.70 0.75 0.47 0.43 0.63
Qwen2.5-7B-Instruct 0.82 0.64 0.68 0.82 0.76 0.76 0.45 0.12 0.58
DeepSeek-Coder-V2-Lite-Instruct 0.79 0.57 0.43 0.82 0.70 0.69 0.41 0.01 0.51
Qwen2.5-Coder-7B-Instruct 0.79 0.64 0.63 0.78 0.78 0.74 0.19 0.12 0.50
Watt-Tool-8B 0.87 0.67 0.54 0.88 0.66 0.76 0.10 0.04 0.47
Hammer2.1-7b 0.76 0.53 0.37 0.60 0.58 0.63 0.26 0.18 0.46
Llama-3.1-8B-Instruct 0.53 0.30 0.28 0.72 0.36 0.45 0.27 0.04 0.34
Phi-3-mini-128k-Instruct 0.48 0.30 0.15 0.58 0.32 0.39 0.25 0.01 0.29
Llama-3.2-3B-Instruct 0.46 0.09 0.09 0.50 0.32 0.33 0.10 0.00 0.22
xLAM-7b-r 0.25 0.02 0.06 0.56 0.00 0.19 0.01 0.07 0.12
Hammer2.1-3b 0.12 0.09 0 0.44 0.08 0.12 0.01 0.01 0.07

### v3

@@ C.2 Special Example
C.3 Agent Example
Figure 20: Example of Agent Data.
Figure 21: Example of Agent Data.
Appendix D Evaluation Inference Prompts
D.1 Normal Prompt
@@ D.3 Agent Prompt
Appendix E Error Examples
E.1 Error Examples of Normal Data
@@
E.2 Error Examples of Special Data
@@
Appendix F Formula for Overall Accuracy
@@
F.1 Detailed Results
Table 6: Accuracy Evaluation of Different Models on English Data (%).
@@ F.1 Detailed Results
GPT-4o 90.0 78.0 68.0 80.0 78.0 82.5 92.7 56.0 81.1
GPT-4-Turbo 90.7 80.5 69.0 80.0 88.0 84.2 82.0 62.5 80.3
Qwen-Max 88.0 75.0 61.0 74.0 82.0 79.7 74.0 60.0 75.1
GPT-4o-Mini 84.3 73.5 59.0 74.0 72.0 76.4 76.7 27.5 68.9
Gemini-1.5-Pro 82.3 73.0 61.0 74.0 72.0 75.7 77.3 26.0 68.5
Claude-3-5-Sonnet 66.7 64.0 46.0 58.0 68.0 62.2 72.7 44.0 62.2
Doubao-Pro-32k 75.3 58.0 52.0 70.0 54.0 66.3 50.7 26.5 56.0
@@
Qwen2.5-Coder-32B-Instruct 86.0 73.5 59.0 76.0 72.0 77.4 80.0 50.0 73.9
DeepSeek-V3 88.0 77.5 63.0 76.0 78.0 80.3 72.7 34.0 71.1
Qwen2.5-72B-Instruct 81.3 74.5 64.0 76.0 80.0 76.8 74.0 37.5 70.0
Llama-3.1-70B-Instruct 83.7 71.5 61.0 74.0 66.0 75.6 29.3 41.0 57.9
Qwen2.5-7B-Instruct 70.3 57.0 49.0 62.0 58.0 62.8 49.3 15.0 51.8
Qwen2.5-Coder-7B-Instruct 73.3 63.5 52.0 70.0 58.0 66.6 25.3 18.5 48.1
DeepSeek-Coder-V2-Lite-Instruct 71.7 58.0 50.0 62.0 60.0 64.0 39.3 2.5 47.9
Watt-Tool-8B 84.7 71.5 57.0 70.0 62.0 74.8 2.0 1.5 44.0
Hammer2.1-7B 71.3 62.5 43.0 64.0 52.0 62.9 3.3 15.0 39.6
Phi-3-Mini-128k-Instruct 66.3 49.0 31.0 58.0 32.0 54.0 12.0 0.0 34.4
MLlama-3.1-8B-Instruct 51.0 49.5 28.0 60.0 56.0 48.1 15.3 6.5 32.9
xLAM-7B-r 61.7 42.0 32.0 66.0 0.0 48.7 4.0 10.0 30.8
Llama-3.2-3B-Instruct 31.7 21.5 9.0 34.0 32.0 26.4 8.7 0.0 17.6
Hammer2.1-3B 32.7 14.0 7.0 36.0 32.0 25.5 0.7 1.5 15.2
Table 7: Accuracy Evaluation of Different Models on Chinese Data (%).
@@
GPT-4o 96.7 91.0 86.0 90.0 88.0 92.7 93.3 71.5 89.6
GPT-4-Turbo 95.7 89.0 86.0 92.0 84.0 91.7 91.3 72.5 88.6
Qwen-Max 94.3 86.0 75.0 92.0 84.0 88.7 74.0 68.5 81.7
GPT-4o-Mini 88.7 78.5 74.0 80.0 84.0 83.4 81.3 39.0 76.0
Claude-3-5-Sonnet 87.0 81.0 79.0 84.0 76.0 83.5 82.0 35.0 75.6
Gemini-1.5-Pro 86.7 80.5 68.0 86.0 84.0 82.2 80.0 25.0 72.8
Doubao-Pro-32k 84.3 53.0 64.0 82.0 78.0 75.0 59.3 23.5 62.8
@@
Qwen2.5-Coder-32B-Instruct 94.3 88.5 83.0 90.0 90.0 90.8 81.3 71.5 85.3
Qwen2.5-72B-Instruct 92.3 86.0 75.0 90.0 82.0 87.3 77.3 52.5 79.3
DeepSeek-V3 95.0 90.5 91.0 90.0 88.0 92.6 73.3 35.0 78.5
Llama-3.1-70B-Instruct 81.3 65.0 66.0 84.0 70.0 75.3 47.3 43.5 62.9
Qwen2.5-7B-Instruct 81.7 63.5 68.0 82.0 76.0 75.9 44.7 12.5 57.8
DeepSeek-Coder-V2-Lite-Instruct 78.7 57.5 43.0 82.0 70.0 68.8 41.3 1.5 51.1
Qwen2.5-Coder-7B-Instruct 78.7 64.0 63.0 78.0 78.0 73.5 19.3 12.5 49.6
Watt-Tool-8B 86.7 67.0 54.0 88.0 66.0 76.3 10.0 4.0 47.4
Hammer2.1-7B 76.0 62.5 37.0 60.0 58.0 62.7 26.0 18.5 46.1
Llama-3.1-8B-Instruct 52.7 30.0 28.0 72.0 36.0 45.0 26.7 4.0 33.8
Phi-3-Mini-128k-Instruct 48.0 29.5 15.0 58.0 32.0 38.9 25.3 1.5 29.5
Llama-3.2-3B-Instruct 45.7 9.0 9.0 50.0 32.0 32.7 10.0 0.0 21.6
xLAM-7B-r 25.3 2.0 6.0 56.0 0.0 18.7 1.3 7.5 12.3
Hammer2.1-3B 12.0 9.0 0.0 44.0 8.0 11.8 1.3 1.5 7.4